OpenStudy (kittiwitti1):

The Value V (in millions of dollars) of a famous painting can be modeled by V = 10e^[kt], where t represents the year, with t = 0 corresponding to 2000. In 2008, the same painting was sold for $65 million. Find the value of the painting in 2014.

OpenStudy (missmob):

isn this physics?

OpenStudy (anonymous):

@kittiwitti1 i guess this is not math....

OpenStudy (kc_kennylau):

Que es?

OpenStudy (kc_kennylau):

Ya veo...

OpenStudy (kc_kennylau):

CuAndo editarAs tU la pregunta?

OpenStudy (kc_kennylau):

cuAl es el problema de hablar en espaNol xD

hartnn (hartnn):

First find the value of 'k' in 2008, t = 8 (2008-2000) and V given = 65 so, 65 = 10e^(8k) find k from here

OpenStudy (kittiwitti1):

@hartnn I used an online calculator and got the answer but my calculator won't get the value (it says Overflow error)? o_o

hartnn (hartnn):

to find the value of k ?? use wolframalpha.com

OpenStudy (anonymous):

@kittiwitti1 have you learned logs?

OpenStudy (anonymous):

ln(6.5)/8 =k

OpenStudy (kittiwitti1):

well I did this:\[65=10e^{kt}\]\[\frac{ln6.5}{8}\approx23.3975\]\[V=10e^{23.3975\times 14}\]and the calculator said ERROR: Overflow. My calculator is a TI-84 Plus graphing, btw

OpenStudy (kittiwitti1):

@thisSucks I'm asking if I got any wrong values because the calculator returned an error message.

hartnn (hartnn):

k = 0.233, not 23.33

hartnn (hartnn):

ln 6.5 /8 = 0.2334

OpenStudy (kittiwitti1):

Oh. My bad o_o I guess online calculators aren't that great.

hartnn (hartnn):

i used google for that :P

OpenStudy (kittiwitti1):

Actually, no I read it wrong, it said "23.---"% so maybe it's right

OpenStudy (kittiwitti1):

So if I fix the decimal it should be right?

hartnn (hartnn):

yes, then its right

OpenStudy (kittiwitti1):

Okay, thanks @hartnn (:

hartnn (hartnn):

thank me when you get correct final answer... and post it too, so that i can verify

OpenStudy (kittiwitti1):

So I did this\[10e^{.233975\times14}\approx264.6042\]I assume the error message was because I put too high an exponent on the first try (Google says the highest exponent the calculator can calculate is 99)?

hartnn (hartnn):

you're correct on both things. welcome ^_^

OpenStudy (kittiwitti1):

Hahah okay, thanks again (:

OpenStudy (kittiwitti1):

@missMob @nonamenever @kc_kennylau please delete comments (: